2. Use the Laplace Transform to solve the initial value problem y-3y+2y=h(t), y(O)=0, y(0)=0, where h (t) = { 0,0<t<4 2, t
Answer #1

15)={2}+74 0<t <4 t>4 Fult alene ta .: We -45 1. 2e-45 -45 ze S S = 2e 4s (585) -45 = e h(t) = 0 = 0+ ult-4) (2-0) ult-4) (2)